Suppose the MPC for south Africa equals 9/10. The tax rate = 1/6 and mY = 1/12 if there is a negative spending shock (decreased net exports) of R300 billion, what will be the effect on total production or income in south africa?
1
Expert's answer
2019-04-16T09:30:20-0400

First we need to find the amount of multiplier:

"m = \frac{1}{1 - c\times(1 - t) + m}" = 1/(1 - 0.9*(1 - 1/6) + 1/12) = 3.

So, a negative spending shock (decreased net exports) of R300 billion will decrease income by 3*300 billion = R900 billion.